the smaller of 2 numbers is three less than the larger. twice the larger plus four times the smaller is 36. find both numbers
Answer by scott8148(6628)

x = y - 3

4x + 2y = 36 ___ 2x + y = 18

substituting ___ 2(y - 3) + y = 18 ___ 3y = 24

solve for y, then substitute back to find x